Bond valuation An investor has two bonds in his portfolio that both have a face value of $1,000 and pay a 12% annual coupon. Bond L matures in 16 years, while Bond S matures in 1 year Assume that only one more interest payment is to be made on Bond S at its maturity and that 16 more payments are to be made on Bond L. a. what will the value of the Bond L be if the going interest rate is 5%? Round your answer to the nearest cent. what will the value of the Bond S be if the going interest rate is 5%? Round your answer to the nearest cent. what will the value of the Bond L be if the going interest rate is 9%? Round your answer to the nearest cent. what will the value of the Bond S be if the going interest rate is 9%? Round your answer to the nearest cent. what will the value of the Bond L be if the going interest rate is 11%? Round your answer to the nearest cent. what will the value of the Bond S be if the going interest rate is 11%? Round your answer to the nearest cent.
Answer #1

Par Value of L = 1000
Maturity = 16
Coupon = 12%*1000 = 120

Par Value of S = 1000
Maturity = 1
Coupon = 12%*1000 = 120

a)Price of L = PV of Coupons(using annuity formula) + PV of Par Value = 120*(1-(1+5%)-16/5% + 1000/(1+5%)16= 1758.64

b) Price of S = PV of Coupons(using annuity formula) + PV of Par Value = 120/(1+5%) + 1000/(1+5%)1= 1066.67

c)Price of L = PV of Coupons(using annuity formula) + PV of Par Value = 120*(1-(1+9%)-16/5% + 1000/(1+9%)16= 1249.38

d) Price of S = PV of Coupons(using annuity formula) + PV of Par Value = 120/(1+9%) + 1000/(1+9%)1= 1027.52

e)Price of L = PV of Coupons(using annuity formula) + PV of Par Value = 120*(1-(1+11%)-16/5% + 1000/(1+11%)16= 1073.79

f) Price of S = PV of Coupons(using annuity formula) + PV of Par Value = 120/(1+11%) + 1000/(1+11%)1= 1009.01

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well